OpenStudy (anonymous):

what is the critical number of 8x ln x

OpenStudy (anonymous):

what does critical number means?

OpenStudy (lgbasallote):

it is a value of x for f'(x)=0 @suroj..as far as i know

OpenStudy (anonymous):

take the derivative (using the product rule) set it equal to zero and solve for x

OpenStudy (anonymous):

d/dx lnx= 1/x knowing this, you can use the product rule 8x(1/x)+8lnx 8(1+lnx)=0 1+lnx=0 lnx=-1 e^-1=x

OpenStudy (anonymous):

\[f'(x)=\frac{8x}{x}+8\ln(x)=8+8\ln(x)\] \[8(1+\ln(x))=0\] \[\ln(x)=-1\] \[x=e^{-1}\]

OpenStudy (anonymous):

what dockworker said

OpenStudy (anonymous):

is this even question of derivative?

OpenStudy (anonymous):

what about the function x^(4/5) (x- 10) it has two critical numbers where A<B what are A and B

OpenStudy (anonymous):

i think you asked this earlier. didn't someone solve it?

OpenStudy (anonymous):

critical number also occurs when the derivative is not defined if this helps you

OpenStudy (anonymous):

ln(x) is not defined when x=0

OpenStudy (anonymous):

i dont remember asking this earlier

OpenStudy (anonymous):

maybe i did. i have a boatload of questions

OpenStudy (anonymous):

\[\frac{d}{dx}x^{\frac{4}{5}}(x-10)=\frac{4}{5}x^{\frac{-1}{5}}(x-10)+x^\frac{4}{5}\] \[=\frac{4}{5x^{\frac{1}{5}}}(x-10)+x^{\frac{4}{5}}\] \[=\frac{4(x-10)+5x^{\frac{1}{5}}(x^{\frac{4}{5}})}{5x^{\frac{1}{5}}}\] \[=\frac{4x-40+5x}{5x^{\frac{1}{5}}}=\frac{9x-40}{5x^{\frac{1}{5}}}\] so the 2 critical numbers are x=40/9 and x=0, assuming i made no mistakes